Required information The Foundational 15 (LO6-1, LO6-3, LO6-4, LO6-5, L06-6, L06-7, LO6-8) (The following information applies Accounting
Answer #1

Selling price per unit = $60,000 / 1,000 = $60

Variable expenses per unit = $39,000 / 1,000 = $39

New variable expenses per unit = $39 + $1 = $40

New fixed expenses = $14,700 + $1,500

= $16,200

New sales in units = 1,000 + 200

= 1,200

New operating income = Sales - Variable expenses - Fixed expenses

= (1,200 * $60) - (1,200 * $40) - $16,200

= $7,800